OpenStudy (anonymous):

solve: log x + log (x+3) = log 10 I am stuck... log (x^2+3x) = log 10 x^2+3x=10 from here i just get weird numbers. helppp

OpenStudy (anonymous):

answer in the back says it should equal 2...

OpenStudy (valpey):

It is true that log(x) + log(x+3) = log(x*(x+3)) \[x^2 +3x-10=0\] \[x^2+5x-2x-10=0\]

OpenStudy (anonymous):

i always forget that whole set one side equal to zero....let me see if i can get it from there

OpenStudy (valpey):

FOIL and the domain of the log function and you got it. Cheers.